OpenStudy (anonymous):

secθ=(13/12) what is cot and sin?

OpenStudy (anonymous):

\[\sec \theta = 13/12\] But \[\cos \theta = 1/ \sec \theta\] Knowing cos, find sin then cot = cos/sin

OpenStudy (anonymous):

For sin, use \[\sin \theta = \sqrt{1- \cos ^{2} \theta }\]